Using a convex lens, f=12 cm and S = 12 cm.  
1
S

  =
1
Expert's answer
2020-10-21T09:52:51-0400
"\frac{1}{f}=\frac{1}{s}+\frac{1}{o}\\\frac{1}{12}=-\frac{1}{12}+\frac{1}{o}\\o=6\ cm"